Shannon Corporation wants to find an equation to estimate some of their monthly operating costs for the operating budget for 2018. The following cost and other data were gathered for 2017​:

Maintenance

Machine

Health

Number of

Shipping

Units

Month

Costs

Hours

Insurance

Employees

Costs

Shipped

January

$4,570

159

$8,620

64

$21,180

7,060

February

$4,565

90

$8,620

71

$24,540

8,180

March

$4,655

237

$8,620

93

$23,760

7,920

April

$4,800

313

$8,620

107

$19,440

6,480

May

$5,160

456

$8,620

88

$17,790

5,930

June

$4,723

290

$8,620

90

$27,600

9,200

July

$4,905

337

$8,620

97

$26,280

8,760

August

$4,921

356

$8,620

86

$25,380

8,460

September

$5,044

427

$8,620

98

$20,220

6,740

October

$5,282

481

$8,620

107

$21,450

7,150

November

$5,665

530

$8,620

96

$16,860

5,620

December

$4,751

274

$8,620

102

$29,040

9,680

1.

Which of the preceding costs is​ variable? Fixed?​ Mixed? Explain.

2.

Using the​ high-low method, determine the cost function for each cost.

3.

Combine the preceding information to get a monthly operating cost function for the Shannon Corporation.

4.

Next​ month, Shannon expects to use 420 machine​ hours, have 120 employees, and ship 9,500 units. Estimate the total operating cost for the month.

Requirement 2. Using the​ high-low method, determine the cost function for each cost.

Begin by selecting the general formula that represents the linear cost function for each cost.

Maintenance

y = a + bX

Health insurance

y = a

Shipping costs

y = bX

Now determine the cost function for each cost. ​(Enter all values in the same format as the general linear formula. If a box is not used in the​ equation, leave the box​ empty; do not enter a zero or select a​ label.)

Cost function

Maintenance

y =

  

+

per machine hour

Health insurance

y =

+

Shipping costs

y =

+

per unit shipped

Requirement 3. Combine the preceding information to get a monthly operating cost function for the Shannon Corporation.

Monthly

operating cost

=

+

per machine hour

+

per unit shipped

Requirement 4. Next​ month, Shannon expects to use 420 machine​ hours, have 120 employees, and ship 9,500 units. Estimate the total operating cost for the month.

Monthly operating cost

= $
Answer #1

1)variable cost varies directly with output thus it remains constant constant in per unit. Fixed cost remains constant in totality irrespective of output .

Mixed cost is a cost that is neither constant in per unit (variable )nor in totality (fixed) .

Maintenance cost

Mixed cost

Health insurance Fixed cost
Shipping cost Variable cost

2)

Maintenance cost :

Highest activity Lowest activity Change [A-B]
Change in cost 5665 4565 1100
Change in activity 530 90 440

y= 4340+ 2.5x where x is number of machine hours

Variable cost per machine hours =change in cost /change in machine hours

                         = 1100 /440

                         = $ 2.5 per machine hour

Highest activity :

Fixed cost = Total cost - total variable cost

                 = 5665 - (530 *2.5)

                  = 5665 - 1325

                  = 4340

can also check at lowest activity >

b)Health insurance (y)= 8620

c)shipping cost(y)=total cost for Jan /total number of units shipped

                      = 21180/7060

                        = $ 3 per unit

y= 0+ 3x

>can also check at any month ,it is $ 3 per unit

3)

Total cost(y) = Total fixed cost +total variable cost

   Y = (4340+8620) + 2.5 per machine hour + 3 per unit shipped

   Y= 12960 + 2.5 per machine hour + 3 per unit shipped

4)Monthly operating cost = 12960 + (2.5*420)+(3*9500)

                          = 12960 +1050 + 28500

                         = $ 42510